Opened 7 years ago

Last modified 7 years ago

#419 assigned flight processing

EM10/06, flight day 265/2011, Italy

Reported by: knpa Owned by: knpa
Priority: alpha 4 medium Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors: adbe

Description (last modified by adbe)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/italy/EM10_06-2011_265_Italy

Data arrived from ARSF via network transfer on 28/09/2011

EM10/06 Scientific objective:
Earthquake study in Italy by examining geology.

EM10/06 Priority: a4m

PI: Ken McCaffrey

Notes: This project was previously flown last year in Greece but since that dataset was confiscated, the project was re-planned for Italy.

Sensors:

Camera (25/11/2011)
Eagle (25/11/2011)
Hawk (25/11/2011)
LiDAR (25/11/2011)

Change History (27)

comment:1 Changed 7 years ago by knpa

  • Description modified (diff)

comment:2 Changed 7 years ago by adbe

Started to process navigation

comment:3 Changed 7 years ago by adbe

Processed navigation

comment:4 Changed 7 years ago by jaho

Starting LiDAR processing.

comment:5 Changed 7 years ago by emca

Blocking LiDAR processing:
Processing shows very low intensity values and max AGC values. This is possibly due to the high altitude of flight during data collection. Querying with Ops.

comment:6 Changed 7 years ago by adbe

Started processing camera

comment:7 follow-up: Changed 7 years ago by adbe

Waiting for rcd logs from Phil (getting them off a disk)

comment:8 Changed 7 years ago by jaho

LiDAR ready for delivery check.

comment:9 Changed 7 years ago by adbe

LiDAR delivery check:

  • Probably should change the data quality remarks, so that it scans better, something like:
    • The intensity values for this flight are very low, it is most likely due to the lines being flown at a high altitude. The intensity values are approximately below 50 and automatic gain control (AGC) values for most points are at a maximum of 255. Thus the elevation values may be incorrect due to the intensity based range correction.
  • There seems to be a slight pitch error in all the lines, don't know how important this is

comment:10 Changed 7 years ago by jaho

Reworded data quality remarks in the ReadMe.

comment:11 Changed 7 years ago by adbe

Processing hyperspectral

comment:12 Changed 7 years ago by adbe

There is an eagle line 14 with no corresponding hawk, it is not mentioned in the logsheet, also the raw is black for all bands and has only 200 lines

comment:13 in reply to: ↑ 7 Changed 7 years ago by adbe

Replying to adbe:

Waiting for rcd logs from Phil (getting them off a disk)

14/10/11

comment:14 Changed 7 years ago by adbe

Hyperspectral ready for delivery check

comment:15 Changed 7 years ago by jaho

Hyperspectral delivery check:

Flawless! Ready to deliver.

comment:16 Changed 7 years ago by adbe

Do not have any rcd logs, will process without

comment:17 Changed 7 years ago by jaho

Reprocessed LiDAR with pitch offset -0.00299919. Ready for delivery check under EM10_06-265-lidar-20111122/.

comment:18 Changed 7 years ago by adbe

Camera ready for delivery check

comment:19 Changed 7 years ago by adbe

LiDAR delivery check:

No problems, ready to deliver

comment:20 Changed 7 years ago by jaho

Camera delivery check:

Not great, but no errors either. Ready to deliver.

comment:21 Changed 7 years ago by adbe

Camera, LiDAR delivered to Ken McCaffrey on 25/11/2011 on disk 104. Hyperspectral delivered to Ken McCaffrey on 25/11/2011 on disk 221

comment:22 Changed 7 years ago by adbe

  • Description modified (diff)

comment:23 Changed 7 years ago by mark1

Hyperspectral XML files have been re-created and re-delivered to PI (via email). This corrects Eagle IFOV value in XML and extra comments included.

comment:24 Changed 7 years ago by adbe

Hyperspectral sct's

lineeaglehawk
1-0.05-0.06
2-0.04-0.04
3-0.09-0.02
40.01-0.03
5-0.02-0.09
60.00-0.07
7-0.04-0.03
8-0.02-0.06
9-0.07-0.15
10-0.01-0.04
11-0.08-0.06
12-0.02-0.05
13-0.09-0.04

comment:25 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from jaho to knpa
  • Status changed from new to assigned

Beginning archiving.

comment:26 Changed 7 years ago by knpa

No rcd kml file/event file due to lack of rcd log files.

comment:27 Changed 7 years ago by knpa

Archived at NEODC.

Moved to our archive disk.

Note: See TracTickets for help on using tickets.